Consider the following dialogue between Gilberto, a student in an introductory economics class, and Gilberto’s teaching assistant, Lucia.

GILBERTO: Hi, Lucia. I was very confused by what the professor was lecturing on today in class. The professor said that in real life many people do not use marginal analysis and mistakenly look at average costs and benefits when making decisions. What did she mean by that?

LUCIA: Hi, Gilberto. A lot of students have had questions about this topic. Let’s go over the concept together.

Marginal analysis involves examining the additional costs and benefits of an activity. Suppose you are deciding how many glasses of wine to drink. Some people may look at the costs and benefits of drinking an entire bottle of wine. Marginal analysis says you should consider the costs and benefits of each glass individually.

Suppose you have drank six glasses of wine. You are considering drinking a seventh glass. If the marginal benefit of the seventh glass is $13 and the marginal cost is $7, then you should   an additional glass.

However, if the marginal benefit of the seventh glass is $3 and the marginal cost is $12, then you should   an additional glass.

GILBERTO: I see! So following this logic, I should continue to drink wine as long as the marginal benefits are   the marginal costs. If the marginal cost of the next glass of wine is greater than the marginal benefit, I should   !

LUCIA: I think you get it now. While this is a very simple example, it does show how marginal analysis can help prevent you from drinking an extra wine and getting sick!
Answer #1

Marginal values are different from average values . Let us assume that a chocolate bar costs $10 . Now average cost will be same as marginal cost in this case . Average cost = Total cost / number of units . So each chocolate bar costs $10 . Now whether we are buying for first time or 10th time , the price will not change . So marginal cost of each chocolate is same as average cost .

But marginal benefit and average benefit will never be same . Let us consider the following example :

Eating the first bar gives us a benefit of $15 ( we consider utility as dollars and not as utils ) . The second bar gives us a benefit of $13 and the third bar gives us a benefit of $8 .

Number of bars Total benefit Average benefit Marginal benefit
1 15 15/1 = 15 15
2 15+ 13 = 28 28/2= 14 28-15 = 13
3 15+13+8=36 36/3 = 12 36-28 = 8

So average benefit when 3rd bar is consumed is still more than cost ( $10 ) but marginal benefit is less than cost ( 8 < 10 ) .

This happens because of the phenomena known as law of diminishing marginal utility . As we go on eating more and more chocolates we become satiated and our utility from further consumption falls .